Bandar Industries Berhad of Malaysia manufactures sporting equipment. One of the companys products, a football helmet for th(For requirements 3 and 4, indicate the effect of each variance by selecting F for favorable, U for unfavorable, and Non
Answer #1

Standard Quantity allowed = 3600*0.66

= 2376

Standard cost allowed= 2376*8

= 19008

Materials cost variance = 19008-20520

= 1512 U

Materials Price Variance = (SP-AP) *AQ

= (8 - 20520/2700)*2700

= 1080 F

Materials Quantity Variance = (SQ-AQ) *SP = (2376-2700)*8

= 2592 U